# thanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ Appendix D Estimate of thermalization timescales
Let us estimate the typical size of the matrix element, $\braket{m}{H_{1}(\lambda_{\text{pre}})}{n}$, which flips $k$ spins. From Eq. (48), we see that $H_{0}(\lambda)$ will converge to its prethermal limit in fRG time of order a few drive cycles $\lambda\sim\Omega^{-1}$. Therefore, we can choose a $\lambda_{\text{pre}}$ so that $\Omega\lambda_{\text{pre}}\gg 1$, but still parametrically shorter than the thermalization time scale, as can be verified a posteriori. Assuming this to be true, we can approximate $U(\lambda)\approx\exp(-\lambda H_{0}(\lambda_{\text{pre}}))$ in Eq. (41). Expanding $H_{1}(\lambda_{\text{pre}})$ in Eq. (41) using the BCH formula,
$H_{1}(\lambda_{\text{pre}})=\sum_{k}\frac{(-1)^{k}\lambda_{\text{pre}}^{k}}{k!}{\mathrm{Ad}}_{H_{0}(\lambda_{\text{pre}})}^{k}H_{1}(0)\,.$ (59)
The commutator with $H_{0}(\lambda_{\text{pre}})$ will increase the range of operators by $\mathcal{O}(1)$. Therefore, terms involving $k$ spin flips are from order $\mathcal{O}(k)$ of the expansion. Since $\lambda_{\text{pre}}\gg\Omega^{-1}$, we bound $\lambda_{\text{pre}}^{k}$ by $\Omega^{-k}$. Using ${\left\|{{\mathrm{Ad}}_{H_{0}(\lambda_{\text{pre}})}^{k}H_{1}(0)}\right\|}\leq J^{k}k!$ [30], the matrix element that involves $k$ spin flips will be bounded by order $(J/\Omega)^{k}$.
@@
Since $\Omega/J\gg 1$, the dependence of $a_{k}$ on $\delta k$ will be weak, and we approximate the summation by $\delta k$ from $\infty$- ∞ to $\infty$. The minimum is reached when the terms of exponent $\delta k$ are balanced by terms of exponent $-\delta k$, meaning that $\exp(J(\lambda-\lambda_{\text{pre}}))\sim\mathcal{O}(1)$. Therefore, we obtain
@@
Appendix E Floquet-Magnus analysis for driven spin chain
We compute the Floquet-Magnus expansion for the spin-chain Hamiltonian, utilizing results derived in Appendix C. The unitary transformation to the rotating frame is
@@ Appendix E Floquet-Magnus analysis for driven spin chain
Appendix F Additional numerical results
  • Effect of transverse field in the $x$ direction: In Eq. (10a), $H_{0}$ is purely classical, and so we analyzed the effect of an additional static transverse field along the $x$ direction, such that
    $H(t)=\sum_{i}\left[-JS_{i}^{z}S_{i+1}^{z}-J_{2}S_{i}^{z}S_{i+2}^{z}+B_{x}S_{i}^{x}+2AS_{i}^{x}\cos(\Omega t)\right]\,.$ (70)
    We have analyzed the norm of $P(\lambda)$ near the dynamical freezing point, just as in the main text. In Fig 10 we show our results for $P(\lambda\to\lambda_{c})$ as a function of $A/\Omega$ near FP1 for different values of $B_{x}$. This demonstrates that the freezing point associated with the first (prethermal) fixed point is essentially independent of $B_{x}$, as evidenced by the minimum of $P(\lambda_{c})$
    Figure 10: $P(\lambda\to\lambda_{c})$ as a function of $A/\Omega$ near FP1 for varying $B_{x}$ for $L=10$. All other parameters are same as in Fig. 3
  • Extent of dynamical freezing over longer fRG time: In the main text, we showed results for the plateau associated with $P(\lambda)$ upto $\lambda\sim 1$. However, the plateau extends until later fRG times, as shown in Fig. 11, both at and away from freezing point for $\Omega=10$.
  • Renormalization of freezing points: To examine the possible renormalization of the freezing point away from the leading-order Floquet-Magnus predictions, we compute numerically the location of the minima of $P(\lambda_{c})$ for various driving frequencies. In Fig. 12, a fine-grid scan of $P(\lambda_{c})$ is performed near FP1 as a function of $A/\Omega$ for two different system sizes ($L=8,~{}10$) using flow ED. We observe that with increasing frequency, the minimum of $P(\lambda_{c})$ approaches the value ($=0.601$) predicted by the zeroth-order Magnus expansion. However, for lower frequencies, there is a slight shift in the location of the freezing point.
    Figure 12: A fine-grid scan for $P(\lambda\to\lambda_{c})$ around FP1, for ($L=8,~{}10$) and two different driving frequencies. The dashed line represents the prediction from the zeroth-order Magnus term.
  • Multiple separated instanton events: In Fig. 13(a), we plot the norm of $H_{1}$ as a function of fRG time using flow ED for a system of size $L=6$. The two well-separated instanton events are clearly visible, where the norms of $H_{1}(\lambda)$ peak at approximately $\lambda\approx 125$ and $\lambda\approx 380$, respectively. In Fig. 13(b), the corresponding norm of $H_{0}(\lambda)$ is shown, displaying a clear step-like feature precisely at the locations where $H_{1}(\lambda)$ peaks. Calculations for longer fRG times are constrained by our numerical precision, which in this case is double precision (binary64).
    Figure 13: (a) Frobenius norm of $H_{1}(\lambda)$ as a function of fRG time at FP1 for $L=6$, $\Omega=1.0J$, $J=1$,$J_{2}=0.9J$. (b) Corresponding Frobenius norm of $H_{0}(\lambda)$. The two instanton events are marked by arrows in (a), along with the corresponding step-like features in (b).

### v2

@@ Appendix E Estimate of thermalization timescales
Let us estimate the typical size of the matrix element, $\braket{m|H_{1}(\lambda_{\text{pre}})|n}$, which flips $k$ spins. From Eq. (51), we see that $H_{0}(\lambda)$ will converge to its prethermal limit in fRG time of order a few drive cycles $\lambda\sim\Omega^{-1}$. Therefore, we can choose a $\lambda_{\text{pre}}$ so that $\Omega\lambda_{\text{pre}}=\zeta$. Here we choose $\zeta\gg 1$ to ensure we are close to the prethermal fixed point, but $\zeta$ is still parametrically smaller than the thermalization time scale, i.e. $\zeta$ is an $\mathcal{O}(1)$ number that does not diverge in the $\Omega/J\to\infty$ limit with $A/\Omega$ kept fixed. This assumption is verified a posteriori by the fact that $H_{0}(\lambda)$ approaches the zeroth order magnus result faster at larger $\Omega$ (see Fig. 4 (b)). Assuming now $H_{0}\approx H_{0}(\lambda_{\text{pre}})$ , we can approximate $U(\lambda)\approx\exp(-\lambda H_{0}(\lambda_{\text{pre}}))$ in Eq. (44). Expanding $H_{1}(\lambda_{\text{pre}})$ in Eq. (44) using the BCH formula,
$H_{1}(\lambda_{\text{pre}})=e^{-\Omega\lambda_{\text{pre}}}\sum_{k}\frac{(-1)^{k}\lambda_{\text{pre}}^{k}}{k!}{\mathrm{Ad}}_{H_{0}(\lambda_{\text{pre}})}^{k}H_{1}(0)\,.$ (62)
The commutator with $H_{0}(\lambda_{\text{pre}})$ will increase the range of operators by $\mathcal{O}(1)$. Therefore, terms involving $k$ spin flips are from order $\mathcal{O}(k)$ of the expansion. We replace $\lambda_{\text{pre}}^{k}$ by $\zeta^{k}\Omega^{-k}$. Using ${\left\|{{\mathrm{Ad}}_{H_{0}(\lambda_{\text{pre}})}^{k}H_{1}(0)}\right\|}\leq J^{k}k!$ [3], the matrix element that involves $k$ spin flips will be bounded by order $e^{-\zeta}(\zeta J/\Omega)^{k}$. In the following analysis, we are interested in the asymptotic limit of $\Omega/J\gg 1$, and we therefore drop the $\zeta$ factor which does not grow with $\Omega/J$.
@@
Since $\Omega/J\gg 1$, the dependence of $a_{k}$ on $\delta k$ will be weak, and we approximate the summation by $\delta k$ from -$\infty$ to $\infty$. The minimum is reached when the terms of exponent $\delta k$ are balanced by terms of exponent $\delta k$, meaning that $(J/\Omega)\exp(J(\lambda-\lambda_{\text{pre}}))\sim\mathcal{O}(1)$. Therefore, we obtain
@@
Appendix F Floquet-Magnus analysis for driven spin chain
We compute the Floquet-Magnus expansion for the spin-chain Hamiltonian, utilizing results derived in Appendix D. The unitary transformation to the rotating frame is